An aluminium can of cylindrical shape contains 500 cm^3 of water. The area of the inner cross section of the can is 125 cm^2. All measurements refer to 10°C. Find the rise in the water level if the temperature increase to 80°C. The coefficient of linear expansion of aluminium= 23x10^-6/°C and average coefficient of volume expansion of water = 3.2x10^-4/°C.

Grade:

2 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Cathy,

Ans:- Let the co-efficient of linear exoansion of aluminium is=a and the coeff of vol expansion of water=y The initial area of the bottom=A0 and initial vol of water is V0

The final area is=At=A0(1+2at)        ( Coeff of lateral expansion= 2 coeff of linear expansion)

The final vol of water=V0(1+yt)

The initial length of water is= 500/125=4cm

the final length is= V0(1+yt)/A0(1+2at)

                     Putting the values we get , the final length=4.0765 cm

Hence the increase in the water level is=(4.0765-4)=.0765 cm(ans)

 

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Cathy !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur

Amrin
6 years ago
∆T=change in temperature=80-10=70℃∆T=V-500/3.2×10^-5×500∆T=V-500/16×10^-270×16×10^-2=V-50011.2=V-500V=511.2=final volume of waterExpansion of aluminium:∆T=A-125/125×23×10^-670=A-125/2875×10^-670×2875×10^-6=A-1250.20125=A-125A=125.20125=final area of aluminium after expansionHeight of water at 10℃=500/125=4Height of water at 80℃=511.2/125.20125=4.083Rise in water level=4.083-4=0.083cm